OpenStudy (anonymous):

evaluae th function for the given value of x f(x)=5^x x=-2

OpenStudy (saifoo.khan):

1/25

OpenStudy (anonymous):

5^(-2)=1/(5^2)=1/25

OpenStudy (anonymous):

f(x)=5^x, x=-2 5^(-2) is the same as 1/(5^2) 1/(5^2)=1/25